### v1

# Pressure2Motion: Hierarchical Motion Synthesis from Ground Pressure with Text Guidance

Abstract
We present Pressure2Motion, a novel motion capture algorithm that synthesizes human motion from a ground pressure sequence and text prompt. It eliminates the need for specialized lighting setups, cameras, or wearable devices, making it suitable for privacy-preserving, low-light, and low-cost motion capture scenarios. Such a task is severely ill-posed due to the indeterminate nature of the pressure signals to full-body motion. To address this issue, we introduce Pressure2Motion, a generative model that leverages pressure features as input and utilizes a text prompt as a high-level guiding constraint. Specifically, our model utilizes a dual-level feature extractor that accurately interprets pressure data, followed by a hierarchical diffusion model that discerns broad-scale movement trajectories and subtle posture adjustments. Both the physical cues gained from the pressure sequence and the semantic guidance derived from descriptive texts are leveraged to guide the motion generation with precision. To the best of our knowledge, Pressure2Motion is a pioneering work in leveraging both pressure data and linguistic priors for motion generation, and the established MPL benchmark is the first benchmark for this task. Experiments show our method generates high-fidelity, physically plausible motions, establishing a new state-of-the-art for this task. The codes and benchmarks will be publicly released upon publication.
Introduction
Motion capture (MoCap) is a pivotal technology in digital animation production and robot control. However, classical motion capture systems are plagued by numerous limitations, such as reliance on wearable devices (in optical MoCap (kirk2004skeletal; guerra2005optical) and inertial MoCap (cloete2008benchmarking; vitali2020determining)), the high cost of equipment (particularly for optical motion capture), and dependence on illumination and visual information (in visual MoCap (mehta2020xnect)). These limitations have raised concerns regarding the cost, efficiency, and visual privacy protection of motion capture. To address this issue, we propose employing ground pressure mats in conjunction with simple textual prompts for motion capture, thereby eliminating the reliance on illumination and wearable devices.
However, synthesizing human motion from ground pressure data is a highly challenging task due to its inherently underdetermined nature. Previous works on human pose prediction based on pressure data have only achieved success in scenarios involving large contact areas, such as predicting the poses of individuals lying in bed (clever2020bodies; wu2024seeing). Regressing human motions in standing poses, such as walking, from ground pressure data presents a far greater challenge. On the other hand, although significant progress has been made in the text-to-motion task (t2m; motiondiffuse), using text prompts as control signals is too unconstrained for a motion capture (MoCap) system. These methods rely on clean, manually specified, and kinematically abstract inputs, such as trajectories or keypoints (omnicontrol; zhong2025sketch2anim), to supply additional information beyond text input. Their architectures are tailored for these well-defined signals and are ill-equipped to handle a control modality that is noisy and physically grounded, such as ground pressure data.
In this paper, we propose Pressure2Motion, a novel generative model that, for the first time, conditions motion synthesis on both high-level semantic text and low-level physical pressure as shown in Figure LABEL:fig:teaser. Our framework introduces a Dual-level Pressure Feature Extractor designed to interpret both the rich semantics of pressure (e.g., identifying which body part is in contact, its direction, and magnitude) and its temporal dynamics. This extractor decodes pressure sequences into two distinct control signals: an overall Pressure-Inferred Movement Trajectory and fine-grained Pressure-Induced Posture Shifts. These signals are then injected into a pre-trained motion diffusion model via a novel Hierarchical Pressure-Modulated Motion Synthesizer, which uses a ControlNet and a parallel Adapter module to ensure the final motion is both plausible and semantically consistent.
To advance this novel research direction, we establish the MPL dataset. Constructed upon the extensive MotionPRO dataset (ren2025motionpro). MPL incorporates fresh, meticulously detailed textual annotations for each motion sequence, thereby establishing the inaugural large-scale, paired (text, pressure, motion) benchmark tailored specifically for this research task. We rigorously evaluate our proposed methodology on this dataset, and comprehensive experimental results demonstrate that Pressure2Motion is capable of generating human motions that are both highly realistic and physically coherent. This achievement heralds a new paradigm in non-visual, privacy-preserving motion capture.
Our contributions can be summarized as follows:
  • To the best of our knowledge, we are the first to formulate and address the novel task of generating full-body human motion from the highly sparse inputs of ground pressure and text prompts.
  • A novel hierarchical generative network is designed to extract multi-level features from the ground pressure signals, which are injected hierarchically into a diffusion model for faithful motion synthesis.
  • By extending a pressure feature injection module on top of a pretrained text-to-motion model, both spatial alignment with ground pressure and semantic plausibility are achieved for the pressure-to-motion task.
  • We introduce the MPL dataset, the first large-scale benchmark with paired text, pressure, and motion data, and demonstrate that our method achieves state-of-the-art performance on the benchmark.
Related Work
Pressure-based Pose Estimation
Figure 2: Samples from our MPL dataset, each pairing ground pressure and motion with five levels of text descriptions.
Traditional pose estimation methods rely on RGB cameras or inertial measurement units (IMUs), which lack reliable foot-ground contact cues and are unsuitable for privacy-sensitive scenarios like healthcare. Early pressure-based methods (yousefi2011bed; davoodnia2021bed; clever2020bodies) used single-frame pressure maps to estimate static postures, but were limited to predefined lying poses and low-resolution settings. Recent methods like PIMesh (wu2024seeing) extend this to short pressure sequences for mesh regression but struggle with dynamic motions like walking. Derived signals like center of pressure (COP)(funk2018learning; tripathi20233d) and foot contact(zhang2024mmvp) offer stability cues but are limited to quasi-static scenarios. Existing pressure datasets are mainly focused on in-bed or insole data, limiting coverage for free-moving activities. Most methods are constrained to short, static actions, with long dynamic sequences remaining a challenge.
Multi-modal approaches, such as BodyPressure (clever2022bodypressure) and PressInPose (gao2024pressinpose), and others (yin2022multimodal; liu2022simultaneously; tandon2024bodymap) combine pressure with RGB, depth, or thermal sensors to improve pose accuracy, but still rely heavily on cameras, compromising privacy. MotionPRO (ren2025motionpro) dataset includes over 12.4M motion frames, significantly expanding coverage compared to earlier datasets like MoYo (tripathi20233d) and PSU-TMM100 (kissos2020beyond). They integrates plantar pressure with RGB-based motion capture, but still treats pressure as an auxiliary signal, underexploring its rich physical semantics. This limits its potential for contact-aware modeling and fails to leverage pressure sensing as a privacy-friendly alternative to vision-based systems.
Controllable Motion Synthesis
Beyond generating human motion from natural language descriptions (guo2020action2motion; petrovich2021action; t2m; petrovich2022temos; jiang2023motiongpt; motiondiffuse; tevet2022human; Mofusion; chen2024text), recent research has increasingly explored controllable motion synthesis, where additional user-defined signals such as spatial constraints are introduced to steer or constrain the synthesis process. There are some approaches (posenetedit; ju2023humansd) that support partial pose constraints, while a prominent line of work focuses on controlling the position of key joints. PriorMDM (shafir2024human) finetunes MDM (tevet2022human) to allow control over end-effector trajectories, while GMD (karunratanakul2023guided) and Trace and Pace (rempe2023trace) guide motion through root joint trajectories. OmniControl (omnicontrol), MotionLCM (dai2024motionlcm) and MaskControl (maskcontrol) generalizes this idea, all leveraging ControlNet (zhang2023adding) for conditioning. Sketch2Anim (zhong2025sketch2anim) further extends this paradigm by leveraging both ControlNet and Adapter (rebuffi2017learning) to generate motion jointly conditioning on action semantics, keyposes and trajectories. While such methods provide flexible and expressive control, they rely on explicit, user-specified signals which are kinematic abstractions lacking physical grounding and cannot capture the continuous, frame-wise nature of real-world contact and balance phenomena.
In contrast, we propose pressure sequences as a dense, physically grounded control signal for motion synthesis. Unlike trajectories or keypoints, pressure maps directly measure the real-world interaction between the body and the ground, capturing contact, force distribution, and center-of-pressure dynamics at each frame, ensuring the generation of realistic and physically plausible motions.
Figure 3: The Pressure2Motion pipeline. We first extract an overall Movement Trajectory $\mathbf{T}_{\text{traj}}$ and fine-grained Posture Shifts $\mathbf{S}_{\text{shift}}$ from pressure maps. These signals are then processed by our adapter branch to provide hierarchical control: a ControlNet encodes the trajectory for global guidance, while an Adapter Block fuses this with posture shifts for local refinement. The resulting features are injected into the pretrained MDM to generate plausible motion aligned with the pressure signals.
Dataset
To explore the role of pressure signals in motion synthesis, we introduce the MPL dataset, including Motion-Pressure-Language multimodal data for pressure-aware motion synthesis. This dataset is constructed by expanding upon the MotionPro dataset, a large-scale motion capture dataset that records ground pressure signals alongside full-body human motion. MPL dataset comprises motion sequences, ground pressure data, RGB video sequences, and corresponding textual descriptions, providing a rich resource for investigating the correspondence between ground pressure signals and human motion. Figure 2 shows samples from our dataset, including various motion sequences, their corresponding ground pressure data, and textual descriptions.
Dataset Statistics
Our MPL dataset consists of motion recordings from 25 subjects with varying heights, weights, and body types. The raw recordings are manually segmented into action-level sequences based on clear semantic boundaries, and then temporally resampled to 20 FPS. This results in a total of 20,944 motion sequences, amounting to approximately 2.3 million frames. The dataset encompasses 400 distinct categories of human motion, spanning everyday activities, traditional fitness routines, aerobic exercises, flexibility movements, and specialized motions for humanoid robotics. Each motion sequence lasts $2\sim 8$ seconds, reflecting diverse temporal dynamics across action types.
In the annotation phase, each RGB image sequence is labeled using Qwen2.5-VL (Qwen2.5-VL), a powerful vision-language model capable of processing and understanding long video sequences. For each motion sequence, we generate five textual descriptions with progressively decreasing levels of detail, resulting in a total of 104,720 descriptions across the dataset. By incorporating multiple descriptive levels, we enrich the variety of textual inputs, thereby enhancing our model’s generalization capability and enabling it to comprehend and synthesize a diverse array of motion styles and variations.
Motion Representation
Unlike MotionPro, which leverages the SMPL (SMPL) parametric human model, our approach employs a more comprehensive motion representation sourced from HumanML3D (t2m), an option that is more aptly tailored for motion synthesis frameworks. This representation includes pelvis velocity, local joint positions, joint velocities, joint rotations (in pelvis space), and binary foot contact indicators. This format ensures fine-grained control over the motion while maintaining compatibility with existing generative models.
Specifically, during data processing, we intentionally exclude any transformations that could alter the global positions of motion joints. These transformations may lead to misalignments between the joints and the pressure data. To ensure proper alignment, we avoid such operations, keeping both the motion representation and pressure signals consistently aligned in global space, thus maintaining the integrity of the pressure-aware motion synthesis process.
Method
Our objective is to synthesize human motion that aligns semantically with a given textual description while ensuring physical consistency with the contact dynamics encoded in the pressure distributions. Given a sequence of pressure maps $\mathbf{P}=\{\mathbf{p}^{i}\}_{i=1}^{N}$ over N frames, where each $\mathbf{p}^{i}\in\mathbb{R}^{H\times W}$ represents a frame-wise pressure map captured by ground pressure sensors, and a corresponding text description $c$, our goal is to synthesize a temporally coherent and physically realistic motion sequence $\mathbf{x}^{1:N}=\{\mathbf{x}^{i}\}_{i=1}^{N}$, where $\mathbf{x}^{i}\in\mathbb{R}^{D}$ represents a single pose and $D$ is the dimension of the representation, reflecting both semantic intent and pressure-based physical dynamics.
In this section, we first outline the basics of the Motion Diffusion Model (MDM), then detail two core components of Pressure2Motion: Dual-level Pressure Semantics Extraction and Hierarchical Pressure-Modulated Motion Synthesis. The complete pipeline is shown in Figure 3.
Preliminary: Motion Diffusion Model
Our model builds upon and expands the capabilities of the Motion Diffusion Model (MDM)  (tevet2022human), which was initially pretrained for generating motions from textual inputs, enabling its adaptation for pressure-sensitive motion synthesis. MDM is a denoising diffusion model that follows the DDPM framework (ho2020denoising). Given a motion sequence $\mathbf{x}$, it assumes $T$ noising steps modeled by the stochastic process:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\sqrt{\alpha_{t}}\mathbf{x}_{t-1},(1-\alpha_{t})I),$} (1)
for a noising step $t\in T$, where $\mathbf{x}_{T}\sim\mathcal{N}(0,I)$. MDM models the denoising process by predicting the clean motion $\hat{\mathbf{x}}_{0}=MDM(\mathbf{x}_{t},t,c)$, given a noised motion $x_{t}$, a noise step $t$, and a textual condition $c$. The model is trained using the standard objective:
$\mathcal{L}_{\text{diff}}=E_{\mathbf{x}_{0}\sim q(\mathbf{x}_{0}|c),t\sim[1,T]}[\|\mathbf{x}_{0}-\hat{\mathbf{x}}_{0}\|_{2}^{2}].$} (2)
New motion sampling from MDM is performed iteratively, following the DDPM framework. At each time step $t$, the clean sample $\hat{\mathbf{x}}_{0}$ is predicted and then noised back to $\mathbf{x}_{t-1}$. This process is repeated from $t=T$ until the final clean motion, $\mathbf{x}_{0}$, is achieved. In our adaptation for pressure-aware motion synthesis, the pressure maps are integrated into the denoising process, providing crucial physical context that guides the generation of temporally consistent and physically realistic motions.
Dual-level Pressure Features Extraction
Extracting meaningful control signals from ground pressure images is essential for guiding the synthesized motion to align with real-world ground contact patterns and physical dynamics. To this end, we adopt a dual-level pressure features extraction strategy that separately models Pressure-Inferred Movement Trajectory and Pressure-Induced Posture Shifts.
The Pressure-Inferred Trajectory is essential for capturing the overall movement path and body alignment during motion. To extract this information, we utilize a feature extraction module $\mathcal{F}_{\text{traj}}$ following the architecture from MotionPro (ren2025motionpro), which includes a ResNet-based (he2016deep) pressure encoder, a temporal information processor, and a fully connected projection layer. Given the sparsity of pressure maps, where valid values are limited and primarily found under the feet during standing, the pressure encoder utilizes a compact ResNet architecture with small convolutional kernels to focus on the localized pressure regions, despite the large size of the pressure map. Temporal dynamics are captured using the temporal information processor, which combines a GRU (chung2014empirical) to model long-term dependencies with a self-attention mechanism to capture short-term correlations in the pressure sequence. The extracted spatial-temporal features are passed through a fully connected layer to generate a compact embedding $\mathbf{T}_{\text{traj}}$, which encapsulates the overall motion trajectory. The extraction of the Pressure-Inferred Trajectory is represented as the encoding of pressure information from each frame:
$\mathbf{T}_{\text{traj}}=\mathcal{F}_{\text{traj}}(\mathbf{P}),\quad\mathbf{T}_{\text{traj}}=\ \{T^{i}\}_{i=1}^{N}.$ (3)
To ensure the model generalizes across various physical conditions, this module is trained separately on a diverse set of pressure-motion pairs augmented with different transformations. Once trained, the module is frozen during the motion synthesizer training phase.
In addition to capturing the overall movement trajectory, pressure maps encode detailed cues for Pressure-Induced Posture Shifts, such as center-of-mass movement, balance transitions, and subtle adjustments in posture. To extract these dynamics, we compute both the raw pressure maps and their temporal differences, which help to capture subtle changes in posture over time, essential for synthesizing realistic and dynamic body movements. These are then combined with a grid-based positional encoding to capture spatial relationships across frames. This combined input is passed through a module $\mathcal{F}_{\text{shift}}$, which consists of a multi-scale convolutional layer followed by a fully connected projection, resulting in a compact representation $\mathbf{S}_{\text{shift}}$ of Pressure-Induced Posture Shifts:
$\mathbf{S}_{\text{shift}}=\mathcal{F}_{\text{shift}}(\mathbf{P},\Delta\mathbf{P},\mathbf{e}),\quad\mathbf{S}_{\text{shift}}=\{S^{i}\}_{i=1}^{N},$ (4)
where $\Delta\mathbf{P}$ represents the temporal difference in pressure maps and $\mathbf{e}$ is the grid-based positional encoding.
The Pressure-Inferred Movement Trajectory offers a high-level guidance for the overall path and body alignment, while the Pressure-Induced Posture Shifts capture the fine-grained, dynamic changes in posture that occur throughout the motion. Together, these features serve as the foundational control inputs, ensuring that the synthesized motion not only reflects the semantic intent expressed in the textual description but also adheres to the physical constraints dictated by the pressure signals.
Hierarchical Pressure-Modulated Motion Synthesis
To leverage both broad-scale trajectories and subtle posture shifts, we introduce an architecture consisting of the pre-trained MDM and the Pressure-Aware Adapter Branch, which serves as the control branch for the MDM.
The pre-trained MDM, $\mathcal{F}_{\theta}$, accepts as input the noisy motion $\mathbf{x}_{t}$, the text description $c$, and the time step $t$. It outputs the predicted clean motion, $\hat{\mathbf{x}}_{0}$, reflecting the semantic alignment with the given text prompt. To enhance the pressure-aware motion synthesis, the Pressure-Aware Adapter Branch is introduced. This branch consists of a ControlNet and multiple Adapter blocks that work in parallel.
The ControlNet module $\mathcal{F}_{\text{Ctrl}}$ is implemented as a trainable variant of the pre-trained MDM, initialized with parameters from the original backbone, with a series of zero-initialized linear layers $\mathcal{Z}$ added to each layer. We then inject the Pressure-Inferred Trajectory embeddings $\mathbf{T}_{\text{traj}}$ directly into the noisy motion sequence $\mathbf{x}_{t}$ through element-wise addition. This modified input is passed through the ControlNet to produce a set of residual features $\mathbf{r}$, which are then added to the corresponding adapter blocks, guiding the synthesis process towards motion that aligns with the Pressure-Inferred Trajectory. The whole process can be written as:
$\mathbf{x}_{t}^{\prime}=\mathbf{x}_{t}+\mathbf{T}_{\text{traj}},\quad\mathbf{r}=\mathcal{F}_{\text{Ctrl}}(\mathbf{x}_{t}^{\prime},t,c).$ (5)
To address the need for leveraging Pressure-Induced Posture Shifts, which reflect subtle and localized dynamics, the Adapter blocks $\mathcal{F}_{\text{Adapt}}$ are introduced. The ControlNet provides high-level guidance by aligning the motion with the overall movement path, while the Adapter blocks refine this motion at a local level, incorporating subtle, pressure-induced posture shifts for more detailed motion synthesis. Operating in parallel with the ControlNet $\mathcal{F}_{\text{Ctrl}}$, the Adapter blocks receive residual features $\mathbf{r}$ from the $\mathcal{F}_{\text{Ctrl}}$ as input, while also integrating Pressure-Induced Posture Shift features $\mathbf{S}_{\text{shift}}$ and text embeddings $c$. This design ensures that the high-level motion trajectory provided by the ControlNet is supplemented with fine-grained adjustments from the Adapter blocks, which captures subtle posture shifts and enhances the overall realism of the synthesized motion. Each Adapter block consists of a self-attention layer, a cross-attention module, and a feed-forward network, producing updated residuals that are injected into the motion denoising network.
This framework enables the synthesis of motion that respects both the dynamics in the pressure signals and the semantic intent encoded in the text descriptions, ensuring hierarchical pressure-aware and semantically consistent motion synthesis. We can get the final predicted clean motion:
$\hat{\mathbf{x}}_{0}^{\prime}=\mathcal{F}_{\theta}(\mathbf{x}_{t},t,c)+\mathbf{r}^{\prime},\quad\mathbf{r}^{\prime}=\mathcal{F}_{\text{Adapt}}(\mathcal{Z}(\mathbf{r}),\mathbf{S}_{\text{shift}},c).$ (6)
Finally, to ensure a strong coupling between the generated motion and the input pressure signal, we introduce a pressure-motion consistency loss alongside the standard diffusion loss. This loss measures the alignment between the synthesized motion and the pressure features. Since pressure primarily reflects foot contact and overall body trajectory, we compute the consistency loss based on the global positions of five key joints: the pelvis root, left/right ankle, and left/right foot. The consistency loss is given by:
$\mathcal{L}_{\text{cons}}(\mathbf{T}_{\text{traj}},\hat{\mathbf{x}}_{0}^{\prime})=\frac{\sum_{n}\sum_{j}\sigma_{nj}\odot\left\lVert\mathcal{E}(\mathbf{T}_{\text{traj}})-R(\hat{\mathbf{x}}_{0}^{\prime})\right\rVert}{\sum_{n}\sum_{j}\sigma_{nj}},$ (7)
where $\sigma_{nj}$ is a binary value indicating whether the control signals contains a value at frame $n$ for joint $j$. The equation for $\sigma_{nj}$ specifies that $\sigma_{nj}=1$ when $j$ is one of the five key joints: pelvis, left ankle, right ankle, left foot, or right foot, and $\sigma_{nj}=0$ otherwise. $\mathcal{E}(\cdot)$ extracts control joints positions and $R(\cdot)$ transforms the motion representation to global absolute joints positions. This consistency loss encourages the generated motion to be aligned with the contact patterns encoded in the input pressure maps. Thus, the total training loss is given by the sum of the diffusion loss and the consistency loss, weighted by factors $\lambda_{\text{diff}}$ and $\lambda_{\text{cons}}$ :
$\mathcal{L}_{\text{total}}=\lambda_{\text{diff}}\mathcal{L}_{\text{diff}}+\lambda_{\text{cons}}\mathcal{L}_{\text{cons}}.$ (8)
Experiments
Experimental Settings
Dataset.
Figure 4: Visual comparisons on the MPL dataset. Yellow denotes the predicted results of different methods; blue represents the ground-truth motions. The motions generated by ours align best with the ground truth, especially in the foot region.
Method FID $\downarrow$ Foot Skating $\downarrow$ LMPJPE $\downarrow$ MPJPE $\downarrow$
Trajectory Error
(>> 50cm) $\downarrow$
R-precision
Top-3 $\uparrow$
Ground Truth 0.002 0.0000 0.0000 0.0000 0.000 0.64
MDM 4.819 0.1029 0.2550 0.2996 0.2744 0.458
MotionDiffuse 3.812 0.1138 0.2305 0.2884 0.2650 0.486
OmniControl 0.315 0.0629 0.1362 0.1719 0.1035 0.523
MaskControl 0.388 0.0617 0.1335 0.1695 0.1009 0.534
Regression 40.015 0.7338 0.4322 0.4896 0.5864 0.127
Ours 0.262 0.0553 0.1273 0.1622 0.1445 0.545
Table 1: Comparison of motion synthesis with pressure control signal on the MPL dataset.
We train and evaluate our approach on the proposed MPL dataset. Data augmentations, including random spatial translations and rotations, are implemented to enhance generalization. The dataset is split into training, validation, and test sets with proportions of 80%, 15%, and 5%. More training details can be found in the Appendix.
Evaluation Metrics.
We follow the evaluation protocol from OmniControl (omnicontrol), combining motion quality from HumanML3D (t2m) and trajectory accuracy from GMD (karunratanakul2023guided). Additionally, we introduce two alignment metrics—Mean Per Joint Position Error (MPJPE) and Lower-body MPJPE (LMPJPE)—to assess the consistency between the synthesized motion and ground pressure signals, with LMPJPE focusing on joints relevant to foot-ground interactions. All evaluations are conducted using a motion evaluator trained on our MPL dataset, following HumanML3D settings.
Comparison Experiments
As no prior work has explored human motion synthesis conditioned on pressure sequences, we reimplement several state-of-the-art motion generation models as baselines, including diffusion-based models (tevet2022human; motiondiffuse; omnicontrol) and masked generative models (maskcontrol). To ensure fair comparisons and enable per-frame pressure-based control, we adapt each baseline accordingly. For MDM and MotionDiffuse, we concatenate the global and local pressure embeddings and append them to the noisy motion input at each denoising step. This allows these models to incorporate pressure signals at each frame, providing a consistent pressure-aware motion synthesis. For OmniControl and MaskControl, we replace the original spatial control inputs with the combined pressure embeddings, enabling these models to condition on pressure in a comparable manner to our approach.
In addition to these baselines, we modify the diffusion process in our method by reducing the multi-step denoising process to a single-step denoising. This adaptation transforms our method into a discriminative regression model. By shifting to a regression-based approach, we aim to investigate the effectiveness of a generative model (which is our approach) in synthesizing motion from sparse pressure data.
The results in Table 1 demonstrate that our method outperforms all baselines across various metrics. We achieve the lowest FID (0.262), indicating our generated motion sequences are closer to real-world data. Our model also significantly reduces foot skating (0.0553), and leads in both MPJPE (0.1622) and LMPJPE (0.1273), showcasing better spatial accuracy, particularly in foot-ground interactions. In terms of R-precision, we outperform baselines with a score of 0.545, demonstrating better motion retrieval performance.
As visualized in Figure 4, our method produces physically realistic motions with accurate foot-ground contact, avoiding the foot-sliding and joint-misalignment artifacts common in baselines.
[…]
TrajErr
(>>50cm) $\downarrow$
w/o MT 0.543 0.0665 0.1943 0.2357 0.3086
w/o PS 0.847 0.0629 0.1555 0.2025 0.2041
w/o CL 0.282 0.0721 0.1550 0.1896 0.1621
Full 0.262 0.0553 0.1273 0.1622 0.1445
Table 2: Ablation study of: Movement Trajectory (MT), Posture Shifts (PS), and the Consistency Loss (CL).
Figure 5: Visualization results of ablation study.
Ablation Study
Our ablation study evaluates each key component, with quantitative results in Table 2 and qualitative examples in Figure 5. Omitting the Pressure-Inferred Movement Trajectory causes significant degradation in alignment-related metrics. Removing Posture Shifts significantly degrades realism (FID), which demonstrates the need for fine-grained adjustments. Furthermore, removing the Consistency Loss increases foot skating, confirming its role in ensuring plausible foot-ground contact. Our complete method outperforms all ablated versions, demonstrating that each component is crucial for synthesizing high-fidelity motion.
Effect of Text on Motion Synthesis
We conducted experiments to evaluate how varying textual inputs, alongside a pressure-only input condition (devoid of any textual content), influence the outcomes of motion synthesis. Figure 6 illustrates the visualization results, where (a) represents the motion generated with only pressure data, while (b), (c), and (d) correspond to motions generated with different textual descriptions.
In all four cases, the model successfully learns from the pressure signals, demonstrating the model’s ability to incorporate the physical dynamics encoded in the pressure data. Additionally, variations in the textual input result in different motion patterns. Specifically, the influence of text is more pronounced on the upper body, while the motion of the lower body remains closely tied to the pressure data. These findings validate the model’s ability to synthesize distinct motions based on varying textual descriptions while still maintaining alignment with the pressure signals.
Figure 6: Comparison of motion synthesis results: (a) generated from pressure only, versus (b-d) generated from pressure with varying text prompts.
Conclusion
In this work, we introduce Pressure2Motion, a pioneering approach for human motion synthesis from ground pressure and text prompts. Eliminating the need for cameras or wearable devices, our method enables privacy-preserving and non-intrusive motion capture. We tackle the ill-posed nature of this task with a hierarchical diffusion model specifically designed to interpret physical signals; its dual-level feature extractor decodes pressure into both broad movement trajectories and fine-grained posture adjustments. Our experiments establish strong baseline performance, demonstrating that Pressure2Motion synthesizes high-fidelity and plausible motion sequences. Furthermore, our introduction of the MPL dataset provides the first benchmark to spur future research in this new direction. Future work may focus on enhancing robustness and expanding the dataset to include more dynamic activities, while also leveraging pressure data in simulated reinforcement learning and exploring its potential for body shape estimation in personalized motion modeling.
Limitations.
Our current work is constrained by the dataset’s limited diversity of motion types. More complex scenarios involving inclined surfaces or real-world interactions have yet to be explored.
Pressure2Motion:
Hierarchical Motion Synthesis from Ground Pressure with Text Guidance
Supplementary Material
1 MPL Dataset Details
(a) Gender Distribution
(b) Height Distribution
(c) Weight Distribution
Figure 7: Distribution of gender, height, and weight among the 25 subjects in the MPL dataset.
The MPL dataset is developed to facilitate research in generating full-body human motion from the highly sparse inputs of ground pressure and text prompts, building on top of the MotionPRO dataset (ren2025motionpro), which contains a large-scale collection of human motion sequences captured using plantar pressure sensors. For our research, we extend this dataset by incorporating textual descriptions.
The raw motion sequences in MotionPro are generally around 10 minutes long and are not segmented based on action semantics. Additionally, actions within each long sequence are often repeated 2-3 times. To address this, we manually segmented the sequences based on clear semantic action boundaries, resulting in 20,944 motion sequences, amounting to approximately 2.3 million frames. Each sequence is temporally resampled to 20 FPS for consistency, lasting $2\sim 8$ seconds, reflecting diverse temporal dynamics across action types.
1.1 Data Distribution
Our MPL dataset comprises motion sequences from 25 subjects with diverse physical characteristics, including a balanced distribution of gender, a wide range of heights and weights, and varying body types. We visualize the distribution of the following attributes:
  • Gender: The dataset includes 12 females and 13 males, with a roughly balanced ratio.
  • Height: The subjects range from 157 $cm$ to 184 $cm$ in height, with an average of 172.1 $cm$, covering both shorter and taller individuals.
  • Weight: The weight distribution spans from 44.05 $kg$ to 108.00 $kg$, with an average of 56.67 $kg$, ensuring the inclusion of both lightweight and heavier subjects.
Figure 7 illustrates the distribution for gender, height, and weight. This coverage enhances the robustness and applicability of our motion synthesis model across real-world variations in body structure.
1.2 Data Processing
We reformatted the SMPL(SMPL) parameters from the MotionPro dataset into a more comprehensive motion representation following the HumanML3D (t2m) convention. Each motion sequence of length $N$ is transformed into a representation of shape $(N,263)$ where each frame encodes the pelvis velocity, local joint positions, joint velocities, joint rotations (in pelvis space), and binary foot contact indicators.
During data processing, we intentionally exclude global operations such as “uniforming skeleton” , “put on floor,” and “rotate to face Z+Z+”, which are part of the default HumanML3D preprocessing pipeline. Uniform skeleton retargeting enforces consistent bone lengths across subjects. While these operations help standardize motion, they may distort the global positions of joints relative to the pressure data. To maintain spatial consistency, we retain the original global coordinates of both motion and pressure, ensuring accurate alignment between the two modalities during synthesis.
One notable characteristic of our processing pipeline is that the first frame of each motion sequence, i.e. the root joint of the initial pose is aligned to the origin of the $XZ$ plane. However, this causes a spatial offset between the motion and the corresponding pressure map in the $XZ$ plane. To account for this, we design a mechanism in pressure feature extraction module to predict and correct this offset using pressure information, ensuring precise spatial alignment for downstream tasks.
1.3 Caption Process
Text descriptions are an integral part of the dataset, providing semantic guidance for the motion generation process. To ensure diversity and semantic richness in textual prompts, descriptions are generated using Qwen2.5-VL (Qwen2-VL), a vision-language model capable of processing long video sequences. Specifically, given a motion clip and a brief action keyword from the original MotionPRO dataset, we provide the RGB video frames and keyword as input to Qwen2.5-VL. The model interprets the human activity within the video context and generates five diverse captions at varying levels of detail. These descriptions range from simple high-level actions (e.g., “The person is walking”) to more intricate and detailed descriptions (e.g., “The person is walking with a slight leftward tilt and right arm movement”).
1.4 Augmentation
Specifically, given a pressure sequence, we apply spatial augmentations (translations and rotations) to simulate real-world variations in global orientation and position. We adjust the global offset of the motion accordingly to maintain spatial alignment with the augmented pressure data. This augmentation helps to simulate real-world variations in body posture and pressure signals.
2 Implementation Details
Our models are implemented in PyTorch and trained on 8 NVIDIA A800 GPUs for a total of 100,000 iterations. We adopt the AdamW optimizer (loshchilov2017decoupled) with a learning rate of $1\times 10^{-5}$. The ControlNet is initialized with pretrained weights from MDM (tevet2022human). During training, the parameters of the Movement Trajectory extraction module and the pretrained MDM are frozen to retain their original representations.
We follow prior works and use CLIP (radford2021learning) to encode text prompts into 512-dimensional embeddings. The output features of both the ControlNet and the Adapter modules are also of size 512 to ensure compatibility with the pretrained MDM architecture. The Pressure-Inferred Movement Trajectory and Pressure-Induced Posture Shifts are extracted with output dimensions of $(B,L,39)$ and $(B,L,256)$ respectively, where $B$ is the batch size and $L=196$ is the sequence length. The 39-dimensional trajectory representation includes the global 3D positions ($XYZ$) of the root, left/right ankles, and left/right toes (total 5 joints × 3 = 15), as well as 6D rotation representations for the left/right ankles and toes (4 joints × 6 = 24).
To improve robustness to text variations, we randomly drop 10% of the text conditions during training. This enables the use of Classifier-Free Guidance (CFG) (ho2022classifier) during inference, where we apply a CFG scale of 5. We adopt a standard DDPM (ho2020denoising) framework with $T=1000$ denoising steps. The control strength $\tau$ for injecting pressure signals is defined as $\tau=\frac{20\hat{\Sigma}_{t}}{L}$, where $\hat{\Sigma}_{t}=\min(\Sigma_{t},0.01)$. We set $\lambda_{\text{diff}}=1$ and $\lambda_{\text{cons}}=5$ throughout all experiments.
3 More Visualization Results
We present more motion synthesis results to further demonstrate the effectiveness of our method in generating human motion from sparse pressure data. Figure 8 showcases a variety of synthesized motions across different scenarios, highlighting the robustness of our approach in the pressure-to-motion task. The first few cases demonstrate typical human actions, such as walking, standing, and some daily activities. Our model generates high-fidelity motions that align well with the pressure data and text prompts, maintaining both physical consistency and semantic plausibility.
Particularly interesting are the last two cases in Figure 8. The second-to-last case corresponds to a jumping motion, where no pressure is applied during the jump. Despite the absence of pressure in the air, our model successfully generates a realistic jumping motion, demonstrating its ability to handle scenarios with no foot-ground contact.
The final case in Figure 8 shows a dynamic plank position, which involves complex pressure distributions from both the hands and feet. Our method effectively handles this rare and specialized pressure contact scenario, generating a physically plausible motion that corresponds to the simultaneous pressure from all four limbs. This demonstrates the versatility of our model in handling uncommon or intricate pressure patterns.
Figure 8: More visualization results on the MPL dataset. Yellow denotes the predicted results of different methods; blue represents the ground-truth motions.
4 Evaluation Details
[…]
TrajErr
(>>50cm) $\downarrow$
w/o ctrlnet 1.3683 0.0621 0.1694 0.1951 0.1650
w/o Adapter 0.695 0.0634 0.1702 0.2092 0.1742
Full 0.262 0.0553 0.1273 0.1622 0.1445
Table 3: Additional Ablation study of Model Archtecture.
We adopt a text feature extractor and a motion feature extractor from HumanML3D and retrain it on our MPL dataset to adapt to the new data distribution. The resulting model is used to evaluate all the methods.
We evaluate motion quality, physical consistency, and semantic alignment of the generated motions using the following metrics:
  • Fréchet Inception Distance (FID) $\downarrow$: FID measures the distributional distance between generated motions and ground-truth motions in the feature space. In our setting, motions are encoded via a pre-trained motion encoder, and FID is computed on the extracted features. Lower FID indicates that the generated motions are more realistic and distributionally similar to real data.
  • Foot Skating $\downarrow$: This metric computes the ratio of frames in which a foot joint is supposed to be in contact with the ground but exhibits non-negligible motion, indicating physically implausible sliding. Specifically, for each frame, we check whether a foot is labeled as ”in contact” and simultaneously has a velocity exceeding a small threshold. The ratio of such inconsistencies over all frames is reported. A lower value indicates better foot-ground contact realism and physical plausibility.
  • Mean Per Joint Position Error (MPJPE) $\downarrow$: MPJPE measures the average Euclidean distance between corresponding joints in the predicted and ground-truth motions:
    $\text{MPJPE}=\frac{1}{N\cdot T}\sum_{t=1}^{T}\sum_{j=1}^{N}\|\hat{\mathbf{p}}_{t,j}-\mathbf{p}_{t,j}\|_{2}$
    where $T$ is the number of frames, $N$ is the number of joints, and $\hat{\mathbf{p}}_{t,j}$ and $\mathbf{p}_{t,j}$ denote the predicted and ground-truth positions of joint $j$ at time $t$. This metric evaluates the spatial alignment between generated and real motions.
  • Lower-body MPJPE (L-MPJPE) $\downarrow$: A variant of MPJPE that only considers lower-body joints (e.g., hips, knees, ankles, feet), which are most relevant to pressure-ground interactions. It reflects the model’s ability to generate physically grounded lower-body motion. Lower is better.
  • Trajectory Error (>> 50cm) $\downarrow$: This metric measures the ratio of motion sequences in which trajectory frames deviate from the ground truth by more than 50 cm. It reflects whether global body movement is consistently aligned with the physical signal.
  • R-Precision (Top-3) $\uparrow$: R-Precision measures the semantic consistency between generated motions and their associated text prompts. We use a joint motion-text encoder to compute the similarity between generated motion features and ground-truth text embeddings. R-Precision@3 reflects whether the correct caption is ranked among the top-3 retrieved results for a generated motion. Higher values indicate better semantic alignment.
5 Additional Ablation Study
In addition to the ablation study on pressure features, we further investigate the impact of model architecture by removing key components, namely the ControlNet and Adapter. Specifically, we modify the architecture by concatenating the two pressure features—Movement Trajectories and Posture Shifts—directly and feeding them into either the ControlNet or Adapter without using the hierarchical structure.
The results of this experiment are shown in Table 3, where we compare the full model with versions that exclude ControlNet and Adapter. For the version without ControlNet, we observe a significant increase in FID (1.3683) and MPJPE (0.1951), indicating that removing the ControlNet impairs the model’s ability to properly align the motion with the pressure signals, resulting in less accurate motion generation. Similarly, removing the Adapter leads to a noticeable degradation in performance, with FID increasing to 0.695 and MPJPE rising to 0.2092. These results demonstrate the critical role of both components in ensuring the high fidelity and physical plausibility of the generated motions. Moreover, concatenating the two pressure features (Movement Trajectories and Posture Shifts) directly and feeding them into either the ControlNet or Adapter results in inferior performance compared to the full model. This suggests that the hierarchical structure, where ControlNet handles the overall movement trajectory and the Adapter fine-tunes the posture shifts, is essential for generating realistic and semantically aligned motions.
Figure 9 provides visual comparisons of the motion sequences generated by the different model variations. These visualizations further confirm that the full model consistently produces motions that are more physically plausible and aligned with the pressure data, especially in areas such as foot-ground contact.
Figure 9: Additional ablation results on the MPL dataset. Yellow denotes the predicted results of different methods; blue represents the ground-truth motions.
6 Limitations
Despite the promising results, our approach is subject to several limitations.
First, the diversity of motion types in the dataset remains relatively limited. While the MPL dataset includes various basic motions such as walking, standing, and sitting, more complex activities—such as interactions on inclined surfaces or with dynamic real-world environments—are not yet covered. Extending the dataset to include more diverse and complex motion types, as well as scenarios involving pressure data from inclined or interacting surfaces (e.g., walking on stairs or engaging with objects), would significantly enhance the robustness and applicability of the model in real-world use cases.
Another major limitation of our model lies in its computational complexity. While our approach demonstrates high fidelity in synthesizing pressure-aware motions, the underlying architecture—specifically the pressure feature extraction module and the hierarchical pressure-modulated motion synthesis framework—is relatively large and computationally demanding. Additionally, the motion generation process, based on the diffusion model, involves multiple denoising steps, making the inference process slower. On a single NVIDIA A800 GPU, generating one motion sequence takes approximately 180 seconds. This extended inference time can become a bottleneck when deploying the model in real-time applications. As a future direction, we plan to investigate strategies, such as more efficient pressure feature extraction techniques, and explore inference optimization to speed up the motion generation process while maintaining high-quality results.